OpenStudy (anonymous):

if a spherical tank 2 m in diameter can be filled with a liquid for $300, find the cost to fill the tanks of a 10 diameter.

OpenStudy (anonymous):

Volume of a sphere is \[(4/3)\pi r ^{3} = (4/3) \pi (d^{3}/8) \] so the relationtip diameter - cost is cubic. So if your diameter is 5 times bigger , cost is 125 times bigger